OpenStudy (anonymous):

Multiply or divide as indicated. x^-8 * x^-2

OpenStudy (anonymous):

x^a * x^b = x^(a+b)

OpenStudy (jdoe0001):

recall that \(\large \bf a^{-n} = \cfrac{1}{a^n}\)

OpenStudy (anonymous):

x^(-8+(-2))=x^(-10)

OpenStudy (anonymous):

so its x^16.......right

OpenStudy (anonymous):

You add the exponents, not multiply them.

OpenStudy (anonymous):

oh ok thanx

OpenStudy (anonymous):

\[x ^{a}*x ^{b}=x ^{a+b}\]